OpenStudy (anonymous):

The first and seventh terms of a sequence are each 10. Starting with the third term, each term is the sum of the previous two terms. What is the fifth term?

OpenStudy (anonymous):

Pleease help me!

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@Juliette32801

OpenStudy (anonymous):

Anyone???

OpenStudy (anonymous):

That one is difficult, and I dont really have an idea.. Im sorry but good luck (:

OpenStudy (anonymous):

Thanks anyways:)

OpenStudy (anonymous):

yw (:

OpenStudy (anonymous):

Anyone have an idea? Hero?

hero (hero):

So basically, what you have is \(10,a_2, a_3, a_4, a_5, a_6, 10\) And starting with the third term, each term is the sum of the previous two terms so: \(10,a_2, (10 + a_2), (2a_2 + 10), (3a_2 + 20), (5a_2 + 30), 10\)

hero (hero):

Now use formula for nth term of a sequence

OpenStudy (anonymous):

That's right....

OpenStudy (anonymous):

Well, I don't think it would be 2a2 +10, it would just be a2 +10

OpenStudy (anonymous):

and it probably isn't + 10 each time, right?

hero (hero):

"Each term is the sum of the previous two terms"

hero (hero):

\(a_2 + a_2 + 10 = 2a_2 + 10\)

hero (hero):

Because \(a_2 + a_2 = 2a_2\)

OpenStudy (anonymous):

but the third term would be a2 +10

hero (hero):

That's what I have bro

hero (hero):

Do you not see?

hero (hero):

10 is the first term \(a_2\) is the second term \(a_2 + 10\) is the third

OpenStudy (anonymous):

you have 2a2 +10

hero (hero):

Look at it again bro.

OpenStudy (anonymous):

so 2 times a2+10

OpenStudy (anonymous):

oh never mind

hero (hero):

If you don't see that I have \(a_2 + 10\) for the third term, I cannot help you.

OpenStudy (anonymous):

I do see

OpenStudy (anonymous):

well the formula is: tn=t1+(n-1)d and I understand the formula, but what's the difference(d)?

OpenStudy (anonymous):

I already said I understand that bro:)

OpenStudy (anonymous):

My problem is the formula: well the formula is: tn=t1+(n-1)d and I understand the formula, but what's the difference(d)?

hero (hero):

You have to solve for it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

let me try

hero (hero):

You're supposed to be brain storming with me bro

hero (hero):

Everything makes sense except for the last term now

hero (hero):

Why did you stop communicating with me? Now is not the time to bail out on me.

hero (hero):

Okay, I have it. Thanks for the help bro.

hero (hero):

We have to solve \(8a_2 + 50 = 10\)

hero (hero):

Since 10 is supposed to be the sum of the previous two terms. We don't need to use the other formula

hero (hero):

So \(a_2 = -5\)

hero (hero):

So the sequence is 10,-5, 5,-.5,5,10

hero (hero):

I'm pretty sure I lost you bro since you are not communicating with me.

hero (hero):

@me1567 are you there?

OpenStudy (anonymous):

oh god I had some computer problems, and I didn't have time to tell you I wasn't there!

hero (hero):

I see. It is finished...

OpenStudy (anonymous):

Glad you were still logged in so I could tell you... I just finished reading all your replies...wow

hero (hero):

Did you understand it?

OpenStudy (anonymous):

yup:) Thanks for explaining it so well...